OpenStudy (anonymous):

An arrow shot horizontally from a cliff at 15 m/s lands 30 m away. How high is the cliff?

OpenStudy (fifciol):

it's really the physics question but let me try:  t=d/v h=gt^2/2 \[h=\frac{ g d^2}{2v^2 }\]